# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ Appendix C Computational Complexity of Algorithms 2 and 4
For simplicity, let us assume that the number of nodes is the same across the templates and the output, that is, assume $N^{s}=N$ for all $s\in[S]$.
Computation of GW Distances: The Gromov–Wasserstein problem is inherently NP-hard due to its formulation as a quadratic assignment problem [31, 50, 10, 47, 74]. In practice, approximation schemes are employed to mitigate this computational burden. For example, if we consider two networks with $N$ and $M$ nodes, respectively, evaluating the GW objective for a fixed coupling matrix incurs a computational cost of $\mathcal{O}(N^{2}M^{2})$ that can be approximated using entropic regularization [46] reducing its cost to $\mathcal{O}((N^{2}M+NM^{2}))$. The original optimization problem can then be approximately solved using a sequence of Sinkhorn iterations, with each iteration cost being dominated by the complexity of evaluating the GW objective. Taking $L$ iterations to convergence, the total cost of solving a single GW problem becomes $\mathcal{O}((N^{2}M+NM^{2})L)$.
In our experiments, we adopt this approach with low entropy regularization, aiming for approximations that closely match the true GW cost and optimal transport plan. More precisely, we use the function ot.gromov.gromov_barycenters from the POT library [21] based on the fixed-point algorithm proposed in [46], which, under low entropy, is similar to Algorithm 1.
@@
    Computation of GW distances: First, one needs to solve (or estimate) as many GW problems as templates ($S$). Thus, extending the discussion above to $S$ templates, the overall complexity to estimate $\{\pi(\mathbf{Y},s)\}_{s\in[S]}$ is $\mathcal{O}((N^{2}M+NM^{2})LS)$, using the approximation through the entropic regularized version of the GW distance.
@@
This yields a total complexity of $\mathcal{O}((N^{2}M+NM^{2})LS+M^{2}S^{2}+S^{3})$ (under the entropic regularization scheme). Since the number of templates $S$ is typically much smaller than the number of nodes $N$, we can conclude that the computational complexity of Algorithm 2 is comparable to solving $S$ Gromov–Wasserstein optimization problems (8). This could be further improved by parallelizing the $S$ Gromov-Wasserstein computations.
@@
    Blow-up step: The computational complexity of Algorithm 3, is essentially equivalent to solving $S$ Gromov-Wasserstein optimization problems. In practice, we approximate each GW computation with its regularized version (using low entropy).
@@
    In the case of classical OT, any optimal coupling–which is a feasible solution of the corresponding linear program–has support size at most $N+M-1$, assuming the source and target discrete measures have support of size $N$ and $M$, respectively (see, e.g., [47, Prop. 3.4]). Linear programming theory does not apply in the GW case due to the quadratic nature of the objective, even though the optimization is performed over the same set of transport plans (or couplings). As a result, no support size bound analogous to that in classical OT is currently known. For empirical insight, we refer the reader e.g. to [14, App., Sec. C], where the authors observe that representations often occur in a space much smaller than the naive requirement of size $NM$.
@@
    Suppose that, after the blow-up step, each $\mathbb{X}^{s}$ has been aligned to $\mathbb{Y}$, so that the collection $\{\mathbb{Y}_{b}=(\mathbf{Y},\mathrm{q}_{b}),\mathbb{X}^{1}_{b}=(\mathbf{X}^{1},\mathrm{q}_{b}),\dots,\mathbb{X}^{S}_{b}=(\mathbf{X}^{S},\mathrm{q}_{b})\}$ consists of networks with $M_{b}$ nodes. Then, similarly to Algorithm 2, computing the matrix $\mathcal{A}$ is of order $\mathcal{O}({M_{b}^{2}S^{2}})$, and solving the quadratic program over the simplex $\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{Q}\lambda$ costs approximately $\mathcal{O}(S^{3})$.
@@ Appendix D Experiments
This section aims to supplement Section 5 with additional visualizations and experiments. App. D.1 and App. D.3 expand on the material in Sections 5.1 and 5.2, respectively. App. D.4 presents an additional application of the proposed GW-BCM.
D.1. Reconstruction Performance of Algorithms 2 and 4 Under the GW-BCM
D.1.1. Experiments with 3D Point Clouds
Figure 11. Three sample point clouds from the category ‘airplane’ in the ModelNet40 dataset. For each point cloud, we randomly select between 300 and 500 samples.
In addition to Figures 6 and 7 in the main text, we include the following:
  • Figure 11: The templates from the ModelNet40 - Princeton 3D Object Dataset [70], used for synthesizing GW barycenters in all the figures dealing with 3D data.
  • Figure 12: By using the templates in Figure 11 and uniform weights $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, in Figure 6 in the main text, we start by synthesizing a GW barycenter following the fixed-point approach (i.e., by using the function ot.gromov.gromov_barycenters from the POT Library [21]). Instead, in the new Figure 12, the synthesis of the initial barycenter is performed as suggested in Proposition 4.3, using a convex combination of some blow-up templates, $\mathbf{Y}=\sum_{s=1}^{S}\mathbf{X}^{s}_{b}$. That is, in Figure 6, the synthetic barycentric shape belongs to a GW barycenter space of the form (10), while in Figure 12, emulates to belong to the space of the form (33).
  • Figure 14: We also include Figure 14, which builds on Figures 7a and 7b: Using the three 3D template point clouds from the ModelNet40 dataset shown in Figure 11, we visualize the performance of the two analysis algorithms (Algorithms 2 and 4) in estimating random GW barycentric coordinates $\lambda$ corresponding to GW barycenters constructed through different techniques.
Figure 12. Reconstructions of synthetic GW-like barycenters built as suggested in Section 3.1, using three (top panel) and two (bottom panel) template point clouds of the same class (‘airplane’), and uniform vector of weights, i.e., $\lambda=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$ (top panel), and $\lambda=(\frac{1}{2},\frac{1}{2})$ (bottom panel). Left-most panels: Unlike Figure 6, here we synthesize convex combinations (top panel) $\mathbf{Y}=\sum_{s=1}^{3}\frac{1}{3}\mathbf{X}^{s}_{b}$ and (bottom panel) $\mathbf{Y}=\sum_{s=1}^{2}\frac{1}{2}\mathbf{X}^{s}_{b}$, where $\mathbf{X}^{s}_{b}$ are blow-up template matrices. According to Remark B.9, in the case of considering 2 templates, such convex combination gives rise to an element in the GW barycenter space generated by those 2 templates. Middle panels: MDS of the matrices given by expression (40) after applying Algorithm 4 to estimate $\lambda$. The computed error between the true $\lambda$-vector and its estimate is numerically zero, within floating-point precision (float64), and the GW distance between the two shapes in the top left and bottom right panels is $GW((\mathbf{Y},\mathrm{q}_{b}),(\mathbf{Y}_{\textrm{barybu}},\mathrm{q}_{b}))\approx 10^{-17}$. Right-most panels: MDS of the matrices given by expression (39) after applying Algorithm 2 to estimate the vector $\lambda$. The computed error between the true $\lambda$-vector and its estimation is of order $10^{-1}$ (top panel) and $10^{-2}$ (bottom panel), and the GW distance between the first and last shapes is $GW((\mathbf{Y},\mathrm{q}_{b}),(\mathbf{Y}_{\textrm{bary}},\mathrm{q}_{b}))\approx 10^{-3}$.
Figure 13. Depiction of 10 randomly generated GW barycenter coordinates $\lambda\in\Delta_{2}$ (blue dots) and their estimations (red crosses) via Algorithms 2 (first column) and 4 (second column). GW barycenters are synthesized for each $\lambda$ using the templates from Figure 11 sampled at the same rate (each with 150 points). Top row: Same as Figure 7a. Bottom row: Templates $\{\mathbf{X}^{s}\}_{s=1}^{S}$ are blown-up to $\{\mathbf{X}_{b}^{s}\}_{s=1}^{S}$ (same size) to allow interpretation of barycenters as convex combinations $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}_{b}^{s}$.
Figure 14. Depiction of 10 randomly generated GW barycenter coordinates $\lambda\in\Delta_{2}$ (blue dots) and their estimations (red crosses) via Algorithms 2 (first column) and 4 (second column). GW barycenters are synthesized for each $\lambda$ using the templates from Figure 11 sampled at different rates, at random, each point cloud having between 50 and 150 points. Top row: Same as Figure 7b. Bottom row: Templates $\{\mathbf{X}^{s}\}_{s=1}^{S}$ are blown-up to $\{\mathbf{X}_{b}^{s}\}_{s=1}^{S}$ (same size) to allow interpretation of barycenters as convex combinations $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}_{b}^{s}$.
Comparison and Analysis of the Experiments Illustrated in Figures 6 and 7 in the Main Text and in the New Figures 12, 13, 14: Algorithm 2 is derived from a fixed-point formulation for computing GW barycenters, compatible with the POT function ot.gromov.gromov_barycenters (Section 3). In contrast, Algorithm 4 is based on defining functional gradients and Karcher means within the GW setting via the blow-up technique (Section 4). The experiments visualized in Figure 6 in the main text and in Figure 12 illustrate the qualitative and quantitative performance of our algorithms in their native settings, while also assessing how well each performs when applied outside its original framework. The results show that Algorithm 2 performs with high accuracy when using fixed-point barycenters (Figure 6, bottom left; Figures 7a and 7b, left; and Figure 14, top left), but less effectively with blow-up-based ones (Figure 12, left-most panel; and Figure 14, bottom left). Conversely, Algorithm 4 performs best when using blow-up-based barycenters (Figure 12, middle panel; and Figure 14, bottom right), but still yields reasonable results with fixed-point ones (Figures 6, 7a, 7b, right panels; and Figure 14, top right). Finally, when the fixed-point iteration scheme is used to synthesize the barycenters, Algorithm 4 for recovering the GW barycentric coordinates performs better when the templates are of the same size (compare the right panels of Figures 7a and 7b).
It is important to highlight that, in the synthesis of point clouds for the experiments in Figure 12 and the bottom panels of Figures 13 and 14, convex combinations of blow-up templates correspond to true GW barycenters when only two templates are involved, as noted in Remark B.9. However, when more than two templates are considered–as in the top panel of Figure 12 and in Figures 13 and 14–such convex combinations do not necessarily yield true GW barycenters. This may help explain why Algorithm 2 performs poorly in the bottom-left panel of Figure14.
D.2. Experiments with 2D Point Clouds
In a similar manner to Figure 6 and Figure 12, Figures 15 and 16 provide a visualization of the performance of Algorithms 2 and 4 on 2D point clouds (instead of 3D) under the GW-BCM assumption. We use the Point Cloud MNIST 2D dataset [23]. On the one hand, in Figure 15 we use the function ot.gromov.gromov_barycenters from the POT Library [21], which is based on a fixed-point algorithm similar to Algorithm 1, to synthesize GW barycenters, and then we apply Algorithm 2 to recover the GW barycentric coordinates $\lambda$. On the other hand, in Figure 16 we synthesize a GW barycenter as a convex combination of some blow-up template matrices (all of the same size); and then we then apply Algorithm 4 to recover the GW barycentric coordinates $\lambda$, that is, the coefficients of the convex combination.
Figure 15. In this figure, we depict the following procedure. First, select a digit class and choose $S$ templates within that class. Then, sample a random vector $\lambda$ from the simplex $\Delta_{S-1}$, and synthesize a GW barycenter using these weights via the function ot.gromov.gromov_barycenters from the POT Library. Next, attempt to recover $\lambda$ via Algorithm 2 and reconstruct the barycenter as in formula (39). MDS is used for visualization. In this example, we consider digits 0, 2, 6, and 8 from the Point Cloud MNIST 2D dataset [23]. We select 3 templates for classes 0 and 2, and 4 templates for classes 6 and 8.
Figure 16. In this figure, we depict the following procedure. First, select a digit class and choose $S$ templates within that class, and perform their blow-up obtaining $S$ matrices $\{\mathbf{X}^{s}_{b}\}_{s=1}^{S}$ of the same size. Then, sample a random vector $\lambda$ from the simplex $\Delta_{S-1}$, and synthesize convex combinations of the blow-up template matrices $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}$. Next, attempt to recover $\lambda$ via Algorithm 4 and reconstruct the barycenter as in formula (40). MDS is used for visualization. In this example, we consider digits 5, 2, 6, and 3 from the Point Cloud MNIST 2D dataset [23]. We select 3 templates for classes 5 and 2, and 5 templates for classes 6 and 3.
D.3. Classification
We provide supplementary material that builds on Section 5.2 (see App. D.3.1), and two additional classification methods based on the proposed GW-BCM approach (see App. D.3.3 and App. D.3.2).
D.3.1. t-SNE in the GW Barycentric Coordinate Space
In addition to Figures 8 and 9, we provide Figures 17, 18, 19 and 20 where, besides showing the t-SNE embedding, we include the K-Means clusters and the corresponding confusion matrix191919 Each cell $(i,j)$ shows the number of samples from true class $i$ that were assigned to cluster $j$, thus a strong diagonal pattern suggests high clustering accuracy, meaning most points were assigned to the correct cluster, while off-diagonal values indicate misclassified samples.. For the first three mentioned figures (17, 18 and 19) we use Algorithm 2 for estimating GW barycentric coordinates, and for the last one (20) we use Algorithm 4.
a Two classes, one template per class.
b t-SNE embedding.
c K-Means clustering.
d Confusion matrix.
Figure 17. As in Figure 8 (A), we select one template from the digit classes 0 and 4, and consider 1800 random training samples from the Point Cloud MNIST 2D dataset [23] corresponding to those classes. To estimate GW barycentric coordinates we applied Algorithm 2. Clustering Accuracy: 0.9 via K-Means.
a Two classes, three templates per class.
b t-SNE embedding.
c K-Means clustering.
d Confusion matrix.
Figure 18. Similarly to Figure 9 (A), we select 3 reference templates from the digit classes 0 and 4, and consider 1800 random training samples from the Point Cloud MNIST 2D dataset [23] corresponding to those classes. To estimate GW barycentric coordinates we applied Algorithm 2. Clustering Accuracy: 0.95 via K-Means.
a Two classes, six templates per class.
b t-SNE embedding.
c K-Means clustering.
d Confusion matrix.
Figure 19. Similarly to Figure 9 (B), we select 6 reference templates from the digit classes 0 and 4, and consider 1800 random training samples from the Point Cloud MNIST 2D dataset [23] corresponding to those classes. To estimate GW barycentric coordinates we applied Algorithm 2. Clustering Accuracy: 0.978 via K-Means. Classes are more separated in comparison to Figures 18 and 18 and thus, the accuracy when clustering is higher
a Two classes, three templates per class.
b t-SNE embedding.
c K-Means clustering.
d Confusion matrix.
Figure 20. Similarly to Figure 18, we select 6 reference templates from the digit classes 0 and 4, and consider 500 random training samples from the Point Cloud MNIST 2D dataset [23] corresponding to those classes. To estimate GW barycentric coordinates we applied Algorithm (3) followed by Algorithm 4. Overall Clustering Accuracy: 0.96 via K-Means.
D.3.2. Minimum GW Barycenter Loss - Experimental Setup
@@
    Predicted class for input ($\text{Predicted class for input }(\mathbf{X},\mathrm{p})=\operatorname*{arg\,min}_{1\leq i\leq n}GW((\mathbf{X},\mathrm{p}),(\mathbf{X}_{\lambda_{i}},\mathrm{p}_{\lambda_{i}})).$ (62)
@@
In conclusion, the performance of this approach is not particularly strong, and since it involves GW synthesis with estimated GW barycentric coordinates, along with multiple GW computations for classification (see (62)), it is computationally expensive. In the next section, we outline an alternative classification method based on a criterion defined directly on the space of GW barycentric coordinates $\lambda$.
D.3.3. Maximum Coordinate
In addition to Figure 3, we provide a complementary illustration that helps interpret the meaning of GW barycentric coordinates. While Figure 3 considers two classes from the Point Cloud MNIST 2D dataset [23], the new Figure 21 extends the setting to three classes.
Description of Figure 21: From the Point Cloud MNIST 2D dataset [23], we consider three digit classes: 0, 4, and 7. We select one reference point cloud per class, and compute and plot GW barycentric coordinates $(\lambda_{1},\lambda_{2},\lambda_{3})$ for each point cloud in the set. The visualization provided in Figure 21 highlights how GW barycentric coordinates capture meaningful structure in the data, with a particularly clear clustering around the template for digit class 0. In contrast, digit classes 4 and 7 exhibit overlap, suggesting the presence of shared geometric shapes between them.
For further classification, the following general approach could be employed:
  1. (1)
    Select templates from each class of the dataset ($k$ samples per class).
  2. (2)
  3. (3)
    Given an input $(\mathbf{X},\mathrm{p})$, compute its GW barycentric coordinates (i.e., apply Algorithm 2 or 4).
  4. (4)
    Classify the input into the class corresponding to the template-label with the highest coordinate in the vector of GW barycentric coordinates.
Indeed, in Figure 3, the predicted labels are assigned by selecting the index corresponding to the highest coordinate in the vector $\lambda$.
D.4. Point Cloud Corruption due to Occlusion

### v2

@@ Appendix C Computational Complexity of Algorithms 2 and 4
For simplicity, let us assume that the number of nodes is the same across the templates, that is, assume $N^{s}=N$ for all $s\in[S]$. Assume the size of the input network $\mathbb{Y}=(\mathbf{Y},\mathrm{q})$ is $M$ (i.e., $\mathbf{Y}]in\mathbb{R}^{M\times M}$ and $\mathrm{q}\in\mathbb{R}^{M}$).
Computation of GW Distances: The Gromov-Wasserstein problem is inherently NP-hard due to its formulation as a quadratic assignment problem [32, 53, 11, 50, 80]. In practice, approximation schemes are employed to mitigate this computational burden. For example, if we consider two networks with $N$ and $M$ nodes, respectively, evaluating the GW objective for a fixed coupling matrix incurs a computational cost of $\mathcal{O}(N^{2}M^{2})$ that can be approximated using entropic regularization [49] reducing its cost to $\mathcal{O}((N^{2}M+NM^{2}))$. The original optimization problem can then be approximately solved using a sequence of Sinkhorn iterations, with each iteration cost being dominated by the complexity of evaluating the GW objective. Taking $L$ iterations to convergence, the total cost of solving a single GW problem becomes $\mathcal{O}((N^{2}M+NM^{2})L)$.
In our experiments, from the POT Library [22], we employed the functions:
  • ot.gromov.gromov_wasserstein, which returns an optimal GW transport plan (by solving the optimization problem via a conditional gradient –Frank-Wolfe– method, for which one iteration requires approximately $N^{2}M+NM^{2}$ operations plus the cost of solving one linear OT subproblem), and
  • ot.gromov.gromov_barycenters (which uses block coordinate descent when optimizing) based on the fixed-point algorithm proposed in [49], which is similar to Algorithm 1.
@@
    Computation of GW distances: First, one needs to solve (or estimate) as many GW problems as templates ($S$). Thus, extending the discussion above to $S$ templates, the overall complexity to estimate $\{\pi(\mathbf{Y},s)\}_{s\in[S]}$ is essentially $\mathcal{O}((N^{2}M+NM^{2})LS)$.
@@
This yields a total complexity of order essentially $\mathcal{O}((N^{2}M+NM^{2})LS+M^{2}S^{2}+S^{3})$. Since the number of templates $S$ is typically much smaller than the number of nodes $N$, we can conclude that the computational complexity of Algorithm 2 is comparable to solving $S$ Gromov-Wasserstein optimization problems (5). This could be further improved by parallelizing the $S$ Gromov-Wasserstein computations.
@@
    Blow-up step: The computational complexity of Algorithm 3, is essentially equivalent to solving $S$ Gromov-Wasserstein optimization problems.
@@
    In the case of classical OT, any optimal coupling–which is a feasible solution of the corresponding linear program–has support size at most N+M1N+M-1, assuming the source and target discrete measures have support of size $N$ and $M$, respectively (see, e.g., [50, Prop. 3.4]). Linear programming theory does not apply in the GW case due to the quadratic nature of the objective, even though the optimization is performed over the same set of transport plans (or couplings). As a result, no support size bound analogous to that in classical OT is currently known. For empirical insight, we refer the reader e.g. to [15, Suppl. Mat., Sec. C], where the authors observe that representations often occur in a space much smaller than the naive requirement of size $NM$.
@@
    Suppose that, after the blow-up step, each $\mathbb{X}^{s}$ has been aligned to $\mathbb{Y}$, so that the collection $\{\mathbb{Y}_{b}=(\mathbf{Y},\mathrm{q}_{b}),\mathbb{X}^{1}_{b}=(\mathbf{X}^{1},\mathrm{q}_{b}),\dots,\mathbb{X}^{S}_{b}=(\mathbf{X}^{S},\mathrm{q}_{b})\}$ consists of networks with $M_{b}$ nodes. Then, similarly to Algorithm 2, computing the matrix $\mathcal{A}$ is of order $\mathcal{O}({M_{b}^{2}S^{2}})$, and solving the quadratic program over the simplex $\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{A}\lambda$ costs approximately $\mathcal{O}(S^{3})$.
@@ Appendix D Experiments
This section aims to supplement Section 5 with additional visualizations and experiments. Suppl. Mat. D.1 expands on the material in Subsection 5.2. Suppl. Mat. D.2 presents an additional application of the proposed GW-BCM.
D.1. Clustering, Classification, and Visualizations
We provide supplementary material that builds on Subsection 5.2.
D.1.1. Additional Visualizations
In addition to Figures 2, we provide a complementary illustration that helps interpret the meaning of GW barycentric coordinates.
Description of Fig. 13: While Figure 2 considers two classes from the Point Cloud MNIST 2D dataset [24], the new Fig. 13 extends the setting to three classes. From the Point Cloud MNIST 2D dataset [24], we consider three digit classes: 0, 4, and 7. We select one reference point cloud per class, and compute and plot GW barycentric coordinates $(\lambda_{1},\lambda_{2},\lambda_{3})$ for each point cloud in the set. The visualization provided in Fig. 13 highlights how GW barycentric coordinates capture meaningful structure in the data, with a particularly clear clustering around the template for digit class 0. In contrast, digit classes 4 and 7 exhibit overlap, suggesting the presence of shared geometric shapes between them.
D.1.2. Additional Classification Experiments
We run additional classification experiments on graph datasets from the public TU Dortmund collection [43], in order to compare our analysis method (in particular, Algorithm 2) with the GW-Dictionary Learning approach proposed in [69]. In both cases, we treat the resulting representations as graph embeddings and perform classification in the corresponding estimated GW-coordinate spaces. The results are reported in Table 2.
For these datasets, pairwise node dissimilarities are no longer Euclidean and are defined as shortest-path distances on the graph. The dimensionality of the resulting (low-dimensional) GW-coordinate spaces was chosen to be the same for both approaches, and is determined as follows: (a) in our method, by analysis with respect to a set of three randomly selected templates per class; (b) in the GW-Dictionary approach, by the unmixing weights associated with a dictionary of $3\times C$ learned atoms, where $C$ denotes the number of classes.
In the GW-Dictionary approach, the dictionary atoms are learned in an offline phase using the full dataset, while the embedding of each graph is obtained by solving a GW linear-like problem to compute the corresponding unmixing weights (please refer to [69] for the precise setup and details). In contrast, our method does not require any training phase to compute the embedding coordinates.
We performed 20 repeated stratified $70/30$ train/test splits. For each split, a 5-fold cross-validation procedure was used to select the optimal number of neighbors $k$ for the $k$-NN classifier, which was then evaluated on the corresponding test split. The embedding times reported in Table 2 correspond to the time required to compute the embedding coordinates for the full dataset. For the GW-Dictionary method (DICT), the reported offline time is the cost of learning the dictionary atoms. The comparable classification accuracies indicate that both embedding strategies capture similar discriminative information from the graph data. While the DICT baseline achieves slightly higher accuracies, our method is competitive (there is not statistically significant difference) and does not require the offline training time reported for learning the dictionary atoms.
Accuracy Avg Embedding Time (s)
Dataset FP DICT FP DICT
PROTEINS $0.6678\pm 0.0249$ $0.6711\pm 0.0269$ 26.93 14.37  (offl. 569.23)
MUTAG $0.6904\pm 0.0558$ $0.7035\pm 0.0557$ 2.29 2.76  (offl. 42.38)
COX2 $0.7656\pm 0.0153$ $0.7741\pm 0.0171$ 10.06 10.77  (offl. 124.10)
Table 2. Classification accuracy and average runtimes. For GW-Dictionary Learning (DICT) [69], embedding time corresponds to unmixing the full dataset once; dictionary learning (atoms) is performed offline and reported in parentheses.
D.1.3. Minimum GW Barycenter Loss - Experimental Setup
The following is an additional classification method based on the proposed GW-BCM approach.
@@ D.1.3. Minimum GW Barycenter Loss - Experimental Setup
    (62) $\text{Predicted class for input }(\mathbf{X},\mathrm{p})=\operatorname*{arg\,min}_{1\leq i\leq n}GW((\mathbf{X},\mathrm{p}),(\mathbf{X}_{\lambda_{i}},\mathrm{p}_{\lambda_{i}})).$
@@
In conclusion, the performance of this approach is not particularly strong, and since it involves GW synthesis with estimated GW barycentric coordinates, along with multiple GW computations for classification (see (62)), it is computationally expensive.
D.2. Point Cloud Corruption due to Occlusion